solve each equation and circle one solution, no solution, or identity
1. 10(t-4)= 5(2t-8)
one solution
identity
no solution

Answer by Edwin McCravy(20054)

10(t-4)= 5(2t-8)

Use the distributive principle:

10t-40 = 10t-40    <--- you can tell from here that it is an
                        identity since both sides are identical.

Add 40 to both sides:

   10t = 10t

Subtract 10t from both sides:

    0t = 0

There are infinitely many solutions because 0t = 0 is true for 
any value of t.  

Edwin